Find the Taylor β€˜s theorem expansion ofΒ 

logz=(z-1)-

(π‘§βˆ’1)

2

2

+

(π‘§βˆ’1)

3

2

βˆ’ β‹―, when|𝑧 βˆ’ 1| < 1.


1
Expert's answer
2021-03-09T05:56:29-0500

We have to find Taylor's expansion of f(z) = log z


"f(z) = log(z) = log(1+z-1)"


"f(1) = log 1 = 0"


"f'(z) = \dfrac{1}{z}" "f'(z) = \dfrac{1}{1} = 1"


"f''(z) = -\dfrac{1}{z^2}" "f''(1) = -1"


"f'''(z) = \dfrac{2\times 1}{z^3}" "f'''(1) = 2"


Hence, by Taylor series


"f(z) = f(a) + f'(z-a).(z-a) + \dfrac{f'''(a)(z-a)^3}{2!} +........."


Hence,


"\implies logz = (z -1) - \dfrac{1}{2}(z -1)^2 + \dfrac{1}{3}(z -1)^3 + ........."